A pharmacist prepared an ointment using 28.5g of zinc oxide instead of the 30.5g required. What is the percent error on the basis of the desired quantity?
Answer #1

Ans) Prepared ointment: 28.5g of zinc oxide

Required ointment: 30.5g of zinc oxide

So 30.5-28.5 = 2g less than required ointment.

Formula: %error = Approx-exact / Exact × 100

= 30.5-28.5/28.5 × 100

= 2/28.5 × 100

= 7.017